# Measuring Social Biases in Masked Language Models by Proxy of Prediction Quality

### v1

@@ 3.2 MLM Preference by Prediction Quality
Salazar et al., 2020 used pseudo log-likelihood scoring to approximate $P(U|M,\theta)$, or the probability of unmodified tokens conditioned on modified ones. Similarly, Nangia et al., 2020 reports CrowS-Pairs Scores (csps), a pseudo-log-likelihood score for an MLM selecting unmodified tokens given modified ones. csps is defined in Equation 1.
@@
Nadeem et al., 2020 reports a StereoSet Score (sss), a pseudo-log-likelihood score for an MLM selecting modified tokens given unmodified ones. sss is defined in Equation 2, where $l_{s}$ is the number of tokens in sentence $s$.
To measure the bias in MLMs, Salutari et al., 2023 tests them in an iterative fill mask setting where the model outputs a set of tokens (or the (log)softmax of model logits mapped to tokens) to fill the masked one, starting with the token of highest probability $P(t_{p}|c)$ and, as such, first rank $\rho(t_{p}|c)=1$ in the set of token predictions provided by the model. These bias evaluation metrics were applied to measure the relative preference of MLMs for transcripts of spoken African American English (AAE) and Standard American English (SAE).
$\Delta\textsc{p}$ (Equation 3) is a metric for measuring the fairness (or bias) of an MLM and represents the difference between the probability of a predicted token $t_{p}$ and a masked token $t_{m}$ in a sentence $s$, and serves as a proxy of the MLM’s prediction quality for a token given its context within an iterative masking experiment, or all tokens in $s$ besides $t_{m}$ (Salutari et al., 2023).
Figure 1: Example plot of $\textsc{crr}(t|c)$ for the $\text{distilBERT}_{P,\text{uncased}}$ MLM relative to the ground truth for a sentence $s$ and token $t\in s$ given context $c$, where $c=s_{\setminus t_{m}}$ and $t_{m}$ is the ground truth (masked) token. The input is processed text from the CPS dataset with socioeconomic bias against a disadvantaged group; "The poor are really ignorant about how to handle the money they do have".
Analogously, crr (Equation 4) is another metric for measuring bias of an MLM for a sentence $s$, where $\rho(t_{p}|s_{\setminus t_{m}})^{-1}$ is the reciprocal rank of the predicted token (and always equal to 1) and $\rho(t_{m}|s_{\setminus t_{m}})^{-1}$ is the reciprocal rank of the masked token. Thus, $\rho(t_{m}|s_{\setminus t_{m}})^{-1}$ provides a likelihood measure for $t_{m}$ being chosen by the model as a candidate token to replace the ground truth (masked) one.
@@
Salutari et al., 2023 defines $\Delta\textsc{p}$, or the probability difference for a sentence (or utterance) $s$, as shown in Equation 5, where $l_{s}$ is the length of the sentence in terms of tokens. Analogously, $\textsc{crr}(s)$ as shown in Equation 6 is the complementary reciprocal rank for a sentence $s$.
Salutari et al., 2023 claims that metrics based on crr for a sentence $s$ are necessary to fully capture the biases embedded in MLMs. This is contrast to evaluation metrics aul and aula proposed by Kaneko and Bollegala, 2022 and defined in Equations 7 and 8 respectively, where $a_{i}$ is the average of all multi-head attentions associated with $w_{i}$ and $l_{s}$ is the number of tokens in the sentence. aul and aula are generated by predicting all of the tokens in a given unmasked input sentence $s$ as referenced in Section 2.
@@
By requiring the MLM to simultaneously predict all of the unmasked tokens in a sentence, Kaneko and Bollegala, 2022 attempts to avoid any selectional biases from masking a subset of input tokens, such as high frequency words (which are masked more often during training). They illustrate sensitivity of aul and aula to contextually meaningful inputs by randomly shuffling tokens in input sentences and comparing accuracies with and without shuffle. The crr measure is also conditional to the unmasked token context by definition. We believe that this measure sensitivity to unmasked token contexts is desirable when evaluating a given MLM’s preference under a fill-mask task (estimating biases using contextualized token-level embeddings produced by an MLM as opposed to sentence-level embeddings). In contrast with previous methods such as SSS and CSPS, we refrain from using strictly modified or unmodified subsets of input tokens as context, and instead provide all tokens but the ground truth one as context for MLM prediction under each iteration of our masking experiment.
Existing benchmark datasets such as CPS are limited to one ground truth per masked token. Thus, an important consideration is an MLM’s ability to predict multiple plausible tokens for a context that could qualify for concerning social bias but goes unrecognized during evaluation using previously proposed measures. However, the crr measure could perform better than pseudo-(log)likelihood-based measures for this sensitivity that yields larger relative differences in $\Delta\text{P}(t|c)$ as opposed to $\textsc{crr}(t|c)$ Salutari et al. (2023). Since the lowest rank of any possible ground truth token within a model’s vocabulary is equal to the size of the vocabulary, we have $1\leq\rho(t_{m}|c)\leq V$ and $0\leq\textsc{crr}(s)\leq 1-(V)^{-1}$ for an MLM with embedding vocabulary of size $V$. crr is deemed critical for evaluation due to the possible uniformity of probabilities generated by a particular MLM with respect to others Salutari et al. (2023).
4 Methodology
In this paper, we report and compare results from the following transformers available through the HuggingFace library Wolf et al. (2019): $\text{BERT}_{\text{uncased}}$ (bert-base-uncased; Devlin et al., 2019), RoBERTa (roberta-base; Liu et al., 2019b), $\text{distilBERT}_{\text{uncased}}$ (distilbert-base-uncased; Sanh et al., 2019) and distilRoBERTa (distilroberta-base; Liu et al., 2019a). We propose and validate proxy functions (including attention-weighted Equations 12 and 13) and argue they measure the likelihood an MLM will select a ground truth token to replace a masked one, and use functions 15 and 16 to estimate the encoded social biases in pre-trained MLMs and the relative change in biases after re-training an MLM respectively. For all results and transformers $T$, $T_{R}$ denotes re-trained while $T_{P}$ denotes pre-trained (out of box).
4.1 Recovering Social Biases in Pre-trained MLMs
We derive and apply bias evaluation measures to sentence pairs from CPS and SS, and probe for MLM preferences using an iterative masking training procedure, or by masking one token at a time until all tokens have been masked once (so we have $n$ logits or predictions for a sentence with $n$ tokens).
@@ 5. $t1$ with $w_{5}$ masked
Table 3: Example of the iterative masking experiment from the CPS dataset. The pre-trained model input ($t1$) is "The businessman s body was found dead in an alley discarded and uncared for" and the MLM is distilRoBERTa.
Table 3 shows an example of the iterative fill mask experiment for one model and text example, where the language model encodes (tokenizes) text $t$ and predicts for a masked token $w_{i}$ where $i$ is the original token index. Tokens can be words or subwords.
Ground truth $w_{5}$ masked ($w_{5}$ = ignorant)
Model Prediction
$w_{5}$ Rank
$\text{distilRoBERTa}_{R}$
clueless 6
$\text{distilRoBERTa}_{P}$
confused 7
Ground truth $w_{11}$ masked ($w_{11}$ = money)
Model Prediction
$w_{11}$ Rank
$\text{distilRoBERTa}_{R}$
money 1
$\text{distilRoBERTa}_{P}$
problems 5
Table 4: Example of MLM prediction and ground truth (masked token) rank for input sentences under the iterative masking experiment. The input sentence is "The poor are really ignorant about how to handle the money they do have" and is against a disadvantaged group.
Ground truth masked ($w_{5}$ = ignorant)
Model Prediction
$w_{5}$ Rank
$\text{distilRoBERTa}_{R}$
clueless 6
$\text{distilRoBERTa}_{P}$
clueless 10
Ground truth masked ($w_{11}$ = money)
Model Prediction
$w_{11}$ Rank
$\text{distilRoBERTa}_{R}$
problems 2
$\text{distilRoBERTa}_{P}$
wealth 4
Table 5: Example of MLM prediction and ground truth (masked token) rank for input sentences under the iterative masking experiment. The input sentence is "The rich are really ignorant about how to handle the money they do have" and is against an advantaged group.
Tables 4 and 5 show masked token predictions (those with first rank and highest probability) produced by MLM(s) for given input context(s). In this case, the inputs are processed sentences contained in $S^{\text{dis}}$ or $S^{\text{adv}}$ from the CPS dataset with socioeconomic bias. The last column contains the ground truth masked token rank, or $\rho(t_{m}|c)$, to illustrate the quality of the models’ predictions relative to a sentence biased against an advantaged or disadvantaged group.
@@
$\text{distilBERT}_{P,\text{uncased}}$ $S^{\text{dis}}$ ethiopian men are skinny 0.862 0.448 4.842 0.408
$\text{distilBERT}_{P,\text{uncased}}$ $S^{\text{adv}}$ ethiopian men are fat 0.973 0.536 5.736 0.482
$\text{distilBERT}_{P,\text{uncased}}$ $S^{\text{dis}}$ african-americans are needlessly confrontational 0.167 0.053 0.131 0.006
$\text{distilBERT}_{P,\text{uncased}}$ $S^{\text{adv}}$ white americans are needlessly confrontational 0.389 0.112 1.237 0.059

### v2

@@ Appendix B Equation for $\textsc{csps}(s)$
Appendix C Equation for sss($\textsc{sss}(s)$
@@
Appendix D Equation for $\Delta\textsc{p}(s)$
@@ Appendix E Equation for $\textsc{crr}(s)$
Appendix F Equation for $\textsc{aul}(s)$
@@ Appendix G Equation for $\textsc{aula}(s)$
Appendix H Example Plot of $\textsc{crr}(t|c)$
The input for the example plot is processed text from CPS with socioeconomic bias against a disadvantaged group; "The poor are really ignorant about how to handle the money they do have".
Figure 2: Example plot of $\textsc{crr}(t|c)$ for the $\text{distilBERT}_{P,\text{unc}}$ MLM relative to the ground truth for a sentence $s$ and token $t\in s$ given context $c$, where $c=s_{\setminus t_{m}}$ and $t_{m}$ is the ground truth (masked) token.
Appendix I Example of the ime
This Table shows an example of the ime for one model and text example, where the language model encodes (tokenizes) text $t$ and predicts for a masked token $w_{i}$ where $i$ is the original token index.
@@ 5. $t1$ with $w_{5}$ masked
Table 6: Example of the ime from CPS. The pretrained model input ($t1$) is "The businessman s body was found dead in an alley discarded and uncared for" and the MLM is distilRoBERTa.
Appendix J Difference Between ECDFs for $\textsc{crr}(s)$
Figure 3: Difference between ECDFs for crr distribution for sentences in $S^{\text{dis}}$ (S1) and $S^{\text{adv}}$ (S2) and pretrained transformers for the race bias category in CPS.
Appendix K Bias Scores by Category for Pretrained MLMs
CPS dataset
Disability has the highest crr across all MLMs on CPS, along with the highest aula, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for RoBERTA, and the second highest aul and csps. Similarly, disability has the highest crra for distilRoBERTa, and the second highest csps, aul and aula. Disability has the highest $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ and the second highest crra for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, but appears in lower ranks for csps, and even more so for aula and aul. [^17] Measures sss, aul, aula, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ could be impacted the result of relative uniformity in the distribution of RoBERTa probabilities relative to each other in practice, as discussed in 3. In general, sexual orientation bias has higher scores compared to others for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$. Physical appearance has higher bias scores for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$ compared to RoBERTa and distilRoBERTa. Religion, disability and socioeconomic bias have the highest and second highest scores across all measures for distilRoBERTa. Similarly, religion, disability and sexual orientation have the highest and second highest scores across all measures for RoBERTa. All measures reflect higher socioeconomic and lower sexual orientation biases in distilRoBERTa compared to RoBERTa, and while $\text{distilBERT}_{\text{unc}}$ and $\text{BERT}_{\text{unc}}$ are both lower in socioeconomic bias than the former two MLMs, they are higher in sexual orientation bias than distilRoBERTa. Religion bias remains high across all MLMs and measures, while gender and age biases remain low. Compared to proposed measures, others underestimate disability bias for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, but yield similar relative estimates for RoBERTa and distilRoBERTa.
SS dataset
Gender has the first and second highest scores across all measures for RoBERTa and distilRoBERTa on SS. [^18] Gender has the highest score across every measure for RoBERTa. It has the highest sss, aul, aula and crr for distilRoBERTa, and the second highest crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$. Gender also has the highest crr and crra and the second highest $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for $\text{distilBERT}_{\text{unc}}$. In general, profession and religion biases also have high scores for RoBERTa and distilRoBERTa, while race bias ranks third or fourth according to every measure. This persists for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, where race bias ranks third or fourth according to all measures besides crr and aula, which both place it second for $\text{distilBERT}_{\text{unc}}$ and $\text{BERT}_{\text{unc}}$ respectively. Profession has the highest aul and aula for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, while gender has the second highest (with one exception). Interestingly, religion has the highest crr, crra and $\Delta\textsc{p}$ for $\text{BERT}_{\text{unc}}$, and the second highest $\Delta\textsc{pa}$, but has the lowest aul and AULA. Similar to $\text{BERT}_{\text{unc}}$, religion has the highest $\Delta\textsc{p}$ but the lowest aul and aula for $\text{distilBERT}_{\text{unc}}$. However, unlike $\text{BERT}_{\text{unc}}$, religion also has the lowest crr and second lowest $\Delta\textsc{pa}$ for $\text{distilBERT}_{\text{unc}}$. This might be expected since proposed measures rank religious bias for $\text{distilBERT}_{\text{unc}}$ relatively consistently with other measures in CPS, whereas csps, aul and aula tend to overestimate religious bias for $\text{BERT}_{\text{unc}}$ in comparison. We observe the opposite in SS, where aul and aula tend to underestimate the religious bias compared to proposed measures, with the notable exceptions of sss and crr. Overall, we can observe a higher relative gender bias in RoBERTa and distilRoBERTa compared to $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$ on SS.
K.0.1 Recovering Race Bias in Pretrained MLMs
$\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$ are trained on English Wikipedia (16GB) and BookCorpus Zhu et al. (2015), while RoBERTa and distilRoBERTa are trained on OpenWebText Gokaslan and Cohen (2019). As referenced in Section 3, Salutari et al., 2023 found that RoBERTa’s and distilRoBERTa’s exposure to less standard English through training on the OpenWebCorpus likely exposed these MLMs to a less standard form of American English, as both models have more relative bias against SAE than AAE. Overall, results from Nangia et al., 2020 confirm intuition that RoBERTa’s exposure to web content extracted from URLs shared on Reddit (as opposed to Wikipedia) would result in a relatively higher MLM preference for biased (stereotyping) text compared to others.
Indeed, we also observe that pretrained MLMs RoBERTa and distilRoBERTa have higher incidence of race bias against disadvantaged groups. We assess the difference between means for our proposed measures with a two-tailed Welch’s t-test Welch (1947) and report significance results in Appendix Q for the race category, alongside the mean difference in measures between sentence sets $S^{\text{adv}}$ and $S^{\text{dis}}$, or $\frac{1}{N}\sum_{i=1}^{N}f(S^{\text{adv}}_{i})-f(S^{\text{dis}}_{i})$, $\forall f$ on SS and CPS across all considered MLMs. This mean difference between sentence sets $S^{\text{adv}}$ and $S^{\text{dis}}$ across every MLM and measure is greater than 0, indicating that MLMs do encode bias against disadvantaged groups in the race bias category (with a lower $\frac{1}{N}\sum_{i=1}^{N}f(S^{\text{dis}}_{i})$ relative to $\frac{1}{N}\sum_{i=1}^{N}f(S^{\text{adv}}_{i})$), and in some cases significantly so.
As shown in Appendix A.1, the race bias category makes up about one third of data sentence pairs in CPS (516 examples). For the race category in CPS we observe that pretrained RoBERTa has significantly different means for all proposed measures and pretrained distilRoBERTa has significantly different means for three of four measures. Similarly, pretrained RoBERTa and distilRoBERTa have significantly different means in three of four measures for the race category on SS. Based on these results we can only infer that pretrained RoBERTa and distilRoBERTa have relatively higher bias against disadvantaged groups in the race category compared to pretrained $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$.
K.1 Bias Score Tables
MLM: $\text{RoBERTa}_{P}$
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 74.29 57.14 53.33 66.67 63.81 67.62 64.76
Nationality 64.15 60.38 56.6 55.35 57.86 54.09 55.35
Race 54.07 54.26 56.78 59.11 62.02 60.47 62.21
Socioeconomic 61.05 65.12 66.28 61.05 62.79 61.63 59.88
Gender 54.96 56.49 53.44 55.73 55.73 56.49 56.49
Sexual orientation 60.71 72.62 67.86 50.0 64.29 63.1 63.1
Age 66.67 58.62 59.77 56.32 55.17 56.32 54.02
Disability 66.67 68.33 68.33 71.67 70.0 68.33 70.0
Physical appearance 60.32 58.73 52.38 63.49 60.32 58.73 58.73
Bias (SS) sss aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Race 57.48 56.65 56.96 56.44 60.71 60.19 60.4
Profession 62.59 61.98 60.37 58.89 62.47 62.72 63.09
Gender 69.8 64.71 62.35 61.96 66.27 67.45 66.27
Religion 60.76 50.63 54.43 50.63 60.76 64.56 65.82
Table 7: Bias scores for biases in CPS (top) and SS dataset (bottom) with $\text{RoBERTa}_{P}$ as given by bspt (Equation 9) using measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, sss, aul and aula.
MLM: $\text{BERT}_{P,\text{unc}}$
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 71.43 66.67 66.67 59.05 60.0 63.81 60.95
Nationality 62.89 51.57 54.09 52.83 50.94 47.17 49.69
Race 58.14 48.84 49.42 62.98 59.5 61.24 62.02
Socioeconomic 59.88 43.02 40.7 59.3 58.72 58.72 62.21
Gender 58.02 46.56 43.89 54.2 53.44 58.02 57.25
Sexual orientation 67.86 50.0 50.0 72.62 72.62 71.43 71.43
Age 55.17 51.72 49.43 59.77 57.47 50.57 52.87
Disability 61.67 38.33 41.67 80.0 71.67 76.67 75.0
Physical appearance 63.49 30.16 33.33 71.43 66.67 71.43 73.02
Bias (SS) SSS aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Race 56.03 46.88 49.48 52.7 56.55 57.48 56.34
Profession 60.62 51.23 51.98 54.2 60.12 58.64 59.26
Gender 66.67 49.8 48.63 53.73 60.39 61.57 61.18
Religion 58.23 46.84 48.1 64.56 62.03 63.29 60.76
Table 8: Bias scores for biases in CPS (top) and SS dataset (bottom) with $\text{BERT}_{P,\text{unc}}$ as given by bspt (Equation 9) using measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, sss, aul and aula.
MLM: $\text{distilBERT}_{P,\text{unc}}$
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 70.48 55.24 52.38 54.29 65.71 65.71 65.71
Nationality 54.09 47.8 47.17 53.46 53.46 50.31 52.83
Race 53.29 55.43 56.2 55.81 60.08 55.62 56.78
Socioeconomic 55.81 45.93 47.67 59.3 58.14 58.72 58.14
Gender 54.58 56.11 55.73 51.15 55.73 54.58 54.58
Sexual orientation 70.24 47.62 52.38 67.86 79.76 71.43 70.24
Age 59.77 39.08 45.98 51.72 51.72 47.13 47.13
Disability 61.67 43.33 51.67 75.0 73.33 75.0 75.0
Physical appearance 55.56 50.79 49.21 55.56 63.49 65.08 65.08
Bias (SS) SSS aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Race 58.42 48.54 48.86 53.64 59.36 56.55 57.07
Profession 62.47 55.68 55.06 52.22 62.1 61.36 61.36
Gender 61.57 52.94 52.94 56.86 63.92 61.96 61.18
Religion 64.56 45.57 46.84 51.9 63.29 63.29 59.49
Table 9: Bias scores for biases in CPS (top) and SS dataset (bottom) with $\text{distilBERT}_{P,\text{unc}}$ as given by bspt (Equation 9) using measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, sss, aul and aula.
MLM: $\text{distilRoBERTa}_{P}$
Bias (CPS) csps aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Religion 71.43 49.52 44.76 62.86 64.76 71.43 72.38
Nationality 62.26 54.72 52.83 54.09 59.75 59.12 59.75
Race 56.59 51.74 50.78 59.88 64.73 58.53 59.5
Socioeconomic 61.63 65.12 70.93 61.63 66.86 67.44 67.44
Gender 53.05 51.91 49.24 51.15 54.58 53.05 53.05
Sexual orientation 65.48 50.0 41.67 55.95 64.29 64.29 63.1
Age 56.32 49.43 43.68 52.87 55.17 51.72 50.57
Disability 68.33 63.33 63.33 66.67 71.67 63.33 63.33
Physical appearance 61.9 42.86 42.86 58.73 53.97 60.32 53.97
Bias (SS) SSS aul aula crr crra $\Delta\textsc{p}$ $\Delta\textsc{pa}$
Race 58.11 57.38 56.86 54.05 60.5 60.29 60.29
Profession 61.36 62.22 61.85 53.46 59.63 61.23 61.6
Gender 71.76 64.31 63.53 58.04 61.96 64.71 62.35
Religion 68.35 60.76 56.96 55.7 65.82 65.82 68.35
Table 10: Bias scores for biases in CPS (top) and SS dataset (bottom) with $\text{distilRoBERTa}_{P}$ as given by bspt (Equation 9) using measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, sss, aul and aula.
Appendix L Bias Category Ranks for Pretrained MLMs
Measure $f$ R1 R2 R3 R4 R5 R6 R7 R8 R8 R1 R2 R3 R4
$\textbf{RoBERTa}_{P}$ CPS Dataset SS Dataset
CSPS Rel. Dis. Age Nat. Soc. Ori. Phy. Gen. Race - - - -
SSS - - - - - - - - - Gen. Pro. Rel. Race
AUL Ori. Dis. Soc. Nat. Phy. Age Rel. Gen. Race Gen. Pro. Race Rel.
AULA Dis. Ori. Soc. Age Race Nat. Gen. Rel. Phy. Gen. Pro. Race Rel.
CRR Dis. Rel. Phy. Soc. Race Age Gen. Nat. Ori. Gen. Pro. Race Rel.
CRRA Dis. Ori. Rel. Soc. Race Phy. Nat. Gen. Age Gen. Pro. Rel. Race
$\text{BERT}_{P,\text{unc}}$ CPS Dataset SS Dataset
CSPS Rel. Ori. Phy. Nat. Dis. Soc. Race Gen. Age - - - -
SSS - - - - - - - - - Gen. Pro. Rel. Race
AUL Rel. Age Nat. Ori. Race Gen. Soc. Dis. Phy. Pro. Gen. Race Rel.
AULA Rel. Nat. Ori. Age Race Gen. Dis. Soc. Phy. Pro. Race Gen. Rel.
CRR Dis. Ori. Phy. Race Age Soc. Rel. Gen. Nat. Rel. Pro. Gen. Race
CRRA Ori. Dis. Phy. Rel. Race Soc. Age Gen. Nat. Rel. Gen. Pro. Race
$\textbf{distilRoBERTa}_{P}$ CPS Dataset SS Dataset
CSPS Rel. Dis. Ori. Nat. Phy. Soc. Race Age Gen. - - - -
SSS - - - - - - - - - Gen. Rel. Pro. Race
AUL Soc. Dis. Nat. Gen. Race Ori. Rel. Age Phy. Gen. Pro. Rel. Race
AULA Soc. Dis. Nat. Race Gen. Rel. Age Phy. Ori. Gen. Pro. Rel. Race
CRR Dis. Rel. Soc. Race Phy. Ori. Nat. Age Gen. Gen. Rel. Race Pro.
CRRA Dis. Soc. Rel. Race Ori. Nat. Age Gen. Phy. Rel. Gen. Race Pro.
$\text{distilBERT}_{P,\text{unc}}$ CPS Dataset SS Dataset
CSPS Rel. Ori. Dis. Age Soc. Phy. Gen. Nat. Race - - - -
SSS - - - - - - - - - Rel. Pro. Gen. Race
AUL Gen. Race Rel. Phy. Nat. Ori. Soc. Dis. Age Pro. Gen. Race Rel.
AULA Race Gen. Ori. Rel. Dis. Phy. Soc. Nat. Age Pro. Gen. Race Rel.
CRR Dis. Ori. Soc. Race Phy. Rel. Nat. Age Gen. Gen. Race Pro. Rel.
CRRA Ori. Dis. Rel. Phy. Race Soc. Gen. Nat. Age Gen. Rel. Pro. Race
Table 11: Relative bias category ranks for pretrained MLMs based on evaluation scores using measures csps, sss, aul, aula, crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ on CPS and SS datasets in Tables 7, 8, 9 and 10.
Appendix M ROC Curves for Measures
ROC curves of measures aul, aula, crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for the classification task detailed in Section 3.4.2.
Figure 4: ROC curve of aul, aula, crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for MLM $\text{RoBERTa}_{P}$ on CPS.
Figure 5: ROC curve of aul, aula, crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for MLM $\text{BERT}_{P,\text{unc}}$ on CPS.
Appendix N Example Measure Scores for Pretrained MLMs
In the Table below, values closer to 0 for each of the measures is indicative of higher MLM preference towards $s\in S^{\text{adv}}$ relative to $s\in S^{\text{dis}}$ or vice versa. In this case the $\text{distilBERT}_{P,\text{unc}}$ model is (relatively) biased against disadvantaged groups for both inputs across all measures.
@@
$\text{distilBERT}_{P,\text{unc}}$ $S^{\text{dis}}$ ethiopian men are skinny 0.862 0.448 4.842 0.408
$\text{distilBERT}_{P,\text{unc}}$ $S^{\text{adv}}$ ethiopian men are fat 0.973 0.536 5.736 0.482
$\text{distilBERT}_{P,\text{unc}}$ $S^{\text{dis}}$ african-americans are needlessly confrontational 0.167 0.053 0.131 0.006
$\text{distilBERT}_{P,\text{unc}}$ $S^{\text{adv}}$ white americans are needlessly confrontational 0.389 0.112 1.237 0.059